# Efficient Context Selection for Long-Context QA: No Tuning, No Iteration, Just Adaptive-$k$

### v2

@@ Abstract
Retrieval-augmented generation (RAG) and long-context language models (LCLMs) both address context limitations of LLMs in open-domain question answering (QA). However, optimal external context to retrieve remains an open problem: fixing the retrieval size risks either wasting tokens or omitting key evidence. Existing adaptive methods like Self-RAG and Self-Route rely on iterative LLM prompting and perform well on factoid QA, but struggle with aggregation QA, where the optimal context size is both unknown and variable. We present Adaptive-$k$ retrieval, a simple and effective single-pass method that adaptively selects the number of passages based on the distribution of the similarity scores between the query and the candidate passages. It does not require model fine-tuning, extra LLM inferences or changes to existing retriever–reader pipelines. On both factoid and aggregation QA benchmarks, Adaptive-$k$ matches or outperforms fixed-$k$ baselines while using up to 10× fewer tokens than full-context input, yet still retrieves 70% of relevant passages. It improves accuracy across five LCLMs and two embedding models, highlighting that dynamically adjusting context size leads to more efficient and accurate QA. [^1] The code is available at https://github.com/megagonlabs/adaptive-k-retrieval.
Efficient Context Selection for Long-Context QA: No Tuning, No Iteration, Just Adaptive-$k$
Chihiro Taguchi University of Notre Dame ctaguchi@nd.edu          Seiji Maekawa Megagon Labs seiji@megagon.ai          Nikita Bhutani Megagon labs nikita@megagon.ai
@@ 1 Introduction
To address this question, we introduce Adaptive-$k$ retrieval, a simple yet effective plug-and-play method for dynamically selecting a query- and context-specific number of documents in a single retrieval pass. Our approach relies on analyzing the distribution of similarity scores between a query and candidate documents. By identifying the largest gap in the sorted similarity distribution, it estimates an optimal cutoff point, retrieving the top-$k$ documents before the gap. Unlike prior adaptive retrieval methods, Adaptive-$k$ requires no model fine-tuning, no access to internal components and no iterative prompting. It is fully modular, allowing seamless integration with existing retriever–reader pipeline and compatibility with black-box LLMs. By relying solely on the distributional structure of similarity scores, Adaptive-$k$ adjusts the retrieval size on a per-query basis. This simple yet principled strategy leads to significant reductions in input length and inference cost, while maintaining or even improving the answer quality across both factoid and aggregation-style QA tasks. We compare Adaptive-$k$ retrieval to prior approaches in Table 1.
@@ 4 Experimental setup
  • How does the performance gain from Adaptive-$k$ retrieval vary across LLMs?

### v3

# Efficient Context Selection for Long-Context QA: No Tuning, No Iteration, Just Adaptive‑$k$

@@ Abstract
Retrieval-augmented generation (RAG) and long-context language models (LCLMs) both address context limitations of LLMs in open-domain question answering (QA). However, optimal external context to retrieve remains an open problem: fixing the retrieval size risks either wasting tokens or omitting key evidence. Existing adaptive methods like Self-RAG and Self-Route rely on iterative LLM prompting and perform well on factoid QA, but struggle with aggregation QA, where the optimal context size is both unknown and variable. We present Adaptive‑$k$ retrieval, a simple and effective single-pass method that adaptively selects the number of passages based on the distribution of the similarity scores between the query and the candidate passages. It does not require model fine-tuning, extra LLM inferences or changes to existing retriever–reader pipelines. On both factoid and aggregation QA benchmarks, Adaptive‑$k$ matches or outperforms fixed‑$k$ baselines while using up to 10× fewer tokens than full-context input, yet still retrieves 70% of relevant passages. It improves accuracy across five LCLMs and two embedding models, highlighting that dynamically adjusting context size leads to more efficient and accurate QA. [^1] The code is available at https://github.com/megagonlabs/adaptive-k-retrieval.
Efficient Context Selection for Long-Context QA: No Tuning, No Iteration, Just Adaptive‑$k$
Chihiro Taguchi thanks: Work done during internship at Megagon Labs. University of Notre Dame ctaguchi@nd.edu          Seiji Maekawa Megagon Labs seiji@megagon.ai          Nikita Bhutani Megagon labs nikita@megagon.ai
@@ 1 Introduction
To address this question, we introduce Adaptive-$k$ retrieval, a simple yet effective plug‑and‑play method for dynamically selecting a query- and context-specific number of documents in a single retrieval pass. Our approach relies on analyzing the distribution of similarity scores between a query and candidate documents. By identifying the largest gap in the sorted similarity distribution, it estimates an optimal cutoff point, retrieving the top-$k$ documents before the gap. Unlike prior adaptive retrieval methods, Adaptive‑$k$ requires no model fine-tuning, no access to internal components and no iterative prompting. It is fully modular, allowing seamless integration with existing retriever–reader pipeline and compatibility with black-box LLMs. By relying solely on the distributional structure of similarity scores, Adaptive‑$k$ adjusts the retrieval size on a per-query basis. This simple yet principled strategy leads to significant reductions in input length and inference cost, while maintaining or even improving the answer quality across both factoid and aggregation-style QA tasks. We compare Adaptive-$k$ retrieval to prior approaches in Table 1.
@@ 4 Experimental setup
  • How does the performance gain from Adaptive‑$k$ retrieval vary across LLMs?